OpenStudy (anonymous):

Working with real numbers. Perform the operations. Reduce all fractions to lowest terms. Pleas show work. 3/4-(-1/2)-5/8

OpenStudy (anonymous):

= 6/8 - (-4/8) - 5/8 = 6/8 + 4/8 - 5/8 = = 5/8

OpenStudy (anonymous):

Thank You.

OpenStudy (anonymous):

welcome